a grocery shopper bought 6 ib of steak and 20 ib of groundbeef the steak cost $5/ib more than the groundbeef and the total bill was $82 what was the price per pound for each type of meat?

Answer by ikleyn(52785)

.
Let B be the price per pound of groundbeef, in dollars.

Then the price per pound for steak is (B+5) dollars.


6 lb of steak cost  6*(B+5) dollars.


20 lb of groundbeef cost  20*B dollars.


An equation for the total is


    6*(B+5) + 20*B = 82    dollars.


Simplify and solve for B :


    6B + 30 + 20B = 82


    26B = 82 - 30 = 52


      B = 52/26 = 2 dollars.


Answer.  The price per pound of groundbeef is $2;  the price per pound of steak is  2+5 = 7 dollars.


Check.  6*7 + 20*2 = 82 dollars.   ! Correct !

Solved.